# CommonPower: Supercharging Machine Learning for Smart Grids

### v1

@@ V Control
V-A Overview
Unlike existing tools, CommonPower enables both the deployment and the training of multi-agent systems with heterogeneous controllers, including safety shields for controller types that do not inherently consider safety constraints. While the user can extend CommonPower with arbitrary controller types, we provide a vanilla implementation of a model predictive controller and an RL-based controller. For training the RL-based controllers, CommonPower offers a common interface for both single-agent and multi-agent training algorithms. This interface realizes the interaction with the underlying power system, including synchronization with any non-RL controllers. We distinguish between a training mode and a deployment mode with fixed parameters where the policies of all RL agents are frozen.
Before going into more detail, we want to demonstrate the effectiveness of our approach with a small example. The pseudo-code in Fig. 4 shows how to use CommonPower to set up a power system consisting of two buildings. One is controlled by an RL agent, the other by an optimal controller. To the best of our knowledge, no existing tool is able to realize such a setup with heterogeneous controllers. With CommonPower, the user only has to specify the safeguarding mechanism and the RL algorithm. The RL agent can be trained with a single line of code. After training, the system with the two controllers can be simulated using the same interface.
@@ V-A Overview
V-B Deployment with Fixed Parameters
We refer to the evaluation of controllers with fixed parameters as deployment. This is not synonymous with deployment on real hardware, in which case RL controllers could keep learning. The control flow for this operation mode is provided by Alg. 1: In each time step $t$, the system loops over all controllers, which compute the control inputs for their respective subsystems. For RL-based controllers, this translates to the execution of the learned policy. Each RL controller has a safety shield that adjusts the predicted action if necessary, as described in Sec. V-D. Subsequently, all control inputs are applied simultaneously to advance the discrete-time system dynamics and transition to the next state.
Algorithm 1 Single run
$t\leftarrow t_{0}$
while $t<t_{0}+T$ do
     $\hat{d}_{[t+1,t+T]}\leftarrow$ forecasts(t)
     $u_{t}\leftarrow$ get_controller_inputs($o_{t}$)
     $c_{t}\leftarrow C(x_{t},d_{t},u_{t})$$\triangleright$ Compute cost
     $x_{t+1}\leftarrow$ additional_updates(t)
     $t\leftarrow t+1$triangleright Step system forward
end while
V-C Training
Besides the loading of pre-trained controllers, we also provide an interface that can be adapted for both single-agent and multi-agent RL training algorithms. For single-agent RL, the standard interface for realizing the interaction with the environment has been established by OpenAI: the gymnasium API (application programming interface) [63].
This API is used by most single-agent RL training algorithms, in particular by all algorithms from the StableBaselines (SB3) repository [64], which are considered to be state-of-the-art. Unfortunately, no similar standard has emerged for multi-agent RL thus far. The multi-agent API PettingZoo [65] provided by OpenAI is tailored to gaming environments with the possibility of agent creation and death during episodes. This would incur unnecessary overhead for the application scenarios we consider since we assume a constant number of active agents. Therefore, we do not adopt the PettingZoo API and instead use an interface similar to the multi-agent environment representation provided by RLLib [66], which is also used by PowerGym [53]. The interface can easily be adapted for other training algorithms using wrapper classes. In particular, we already provide a wrapper to adapt it to the gymnasium API for single-agent RL as well as the multi-agent proximal policy optimization (MAPPO) implementation by [57].
V-D Safeguarding RL-based Controllers
TABLE III: Comparison of the realized and potential cost for three time periods of 10 days for optimal (OC) and RL control. The mean and standard deviation of the cost over five training seeds are provided for RL controllers.
Decentralized scenario Centralized scenario
OC RL OC noisy RL noisy OC RL OC noisy RL noisy
29.03.2016 - 07.04.2016 $42.90$ $48.9\pm 0.09$ $43.94$ $48.34\pm 0.07$ $42.90$ $48.88\pm 0.59$ $43.94$ $48.0\pm 0.90$
09.06.2016 - 18.06.2016 $-9.97$ $4.57\pm 0.27$ $-8.52$ $3.93\pm 0.35$ $-9.97$ $5.95\pm 1.96$ $-8.99$ $3.90\pm 1.51$
13.12.2016 - 22.12.2016 $283.25$ $294.72\pm 0.21$ $284.89$ $293.97\pm 0.24$ $283.25$ $294.94\pm 0.08$ $284.67$ $293.63\pm 0.61$
Figure 5: Training curves for the decentralized (MAPPO) and centralized (PPO) scenario. For each training algorithm, we compute the mean and standard deviation over five random seeds of the number of action corrections (left) and the constraint violation penalty (right).
@@ V-D Safeguarding RL-based Controllers
To train and deploy RL-based controllers in safety-critical real-world applications, such as power systems, constraints of the controlled system have to be fulfilled at all times. In the case of decentral control with shared constraints, we adopt a two-stage approach: First, each agent computes its dispatch, and then the grid operator computes the dispatch of balancing assets to fulfill global constraints. This is in line with existing work [42, 43, 35, 45, 44]. In contrast, a single-stage distributed optimization with global constraints requires communication between agents. While we plan to include this in a future release of the CommonPower, we have initially decided on the two-stage approach since it has significantly lower computational overhead and, in our view, is a more realistic representation of real-world power grids in the mid-term future. Furthermore, it is privacy-preserving since each coalition can compute its own dispatch subject to local constraints based on purely local observations, and the grid operator only needs information about the total energy consumption.
@@
The safety shield of an individual agent solves an optimal control problem that captures the constraints of the subsystem under control over a horizon $T$. The output of the safety layer, the safe action for the current time step $t$, is passed to the environment as shown in Fig. 1. The unique feature of CommonPower is that the constraints of the subsystem are automatically extracted from the symbolic domain of the respective entities as described in Sec. IV-B.
The default safeguarding mechanism in CommonPower is based on the predictive safety filter from [20] and formulated as action projection computing a safe action $\tilde{u}^{i}_{t}$ minimizing the Euclidian distance to the proposed action $u^{i}_{t}$ of the RL agent. Accordingly, the safety layer solves
[…]
@@
such that the tuple used for learning is $(o^{i}_{t},u^{i}_{t},o^{i}_{t+1},\tilde{r}^{i}_{t})$. This penalty informs the agent about the intervention of the safety layer. How to compute the penalty is a design choice and can require careful tuning.
VI Experiment
Figure 6: Comparison of the energy management of an optimal controller and an RL controller for a given household with a battery. Fig. 6a compares an RL controller trained just for this household (decentral control) to an optimal controller. In the centralized scenario, the RL controller was trained to control the entire system. Fig. 6b shows the results for the same household as in 6a. We compare deployment under perfect and noisy forecasts. The red crosses indicate where actions had to be corrected.
To comprehensively demonstrate the use of CommonPower, we conduct an experiment comparing RL to an optimal control baseline for a centralized and a decentralized control scheme under consideration of noisy forecasts. The presented case study is an example of a comparison along multiple dimensions, which could not be studied with any existing tool (see Tab. I).
@@ VI-A System Model
The topology of our test system is based on the Kerber Landnetz Kabel 2 network taken from the pandapower library. This network is based on a theoretical categorization of power networks and represents an average rural network [67]. The network has 30 buses, 14 of which are households, and is connected to the external grid via a substation. The topology is imported from pandapower and all parameters are maintained, e.g., line admittances. We use the DC power flow model [68, Ch. 6.2.4]. During import, components are added to households in a stochastic fashion. Each household has an inflexible load with a probability of 100%, a battery with a probability of 50%, and a photovoltaic generator with a probability of 50%. All component parameters are independently and uniformly chosen from a specified range. The generation/load profiles are based on the SimBench [69] dataset 1-LV-rural2–1-sw [^2] https://simbench.de/de/datensaetze. The resulting system is used for training and deployment across all control schemes.
VI-B Control Schemes
In the centralized scenario, we model an energy community in which all generation and storage assets are shared and their dispatch is controlled by a central entity. We model the substation as a trading node at which the central controller can buy and sell energy. To model realistic electricity prices, we use spot market prices from Denmark [^3] https://www.energidataservice.dk/tso-electricity/Elspotprices. We do not make any assumptions on cost sharing within the microgrid and only evaluate the total cost.
In the decentralized scenario, every household individually controls its assets, optimizing its energy cost. Even though CommonPower enables heterogeneous control schemes with both RL and optimal controllers, we intentionally focus on homogeneous control scenarios for clarity. For every household, we assume electricity prices to be identical to the prices of the trading node in the centralized scenario. As a metric for comparison, we use the sum of all household costs.
VI-C RL Training
For the decentralized scenario, each household with a controllable component – in this case a battery – is assigned an RL controller. All RL agents are then trained simultaneously using MAPPO [57]. The observation and action space of a household depend on its setup: A household $i$ with a battery, a photovoltaic generator, and an inflexible load observes
$o_{t}^{i}~{}=~{}\begin{bmatrix}soc_{t}^{i}&\hat{\ell}_{[t,T]}^{i}&\hat{g}_{[t,T]}^{i}&\hat{\phi}_{[t,T]}\end{bmatrix}^{T}$
@@ VI-C RL Training
where $soc_{t}^{i}$ is the state of charge of the battery and ^$\hat{\ell}_{[t,T]}^{i}$, $\hat{g}_{[t,T]}^{i}$, and $\hat{\phi}_{[t,T]}$ are the forecasted load, generation, and electricity prices. The action of such a household is the charging/discharging power of the battery, $u_{t}^{i}=\begin{bmatrix}p_{t}^{B}\end{bmatrix}.$ The state of charge of each battery evolves according to a simplified version of the equation given in Tab. II, with self-discharge $\eta_{s}=1$ and a unified charge/discharge efficiency $\eta_{c}~{}=~{}\frac{1}{\eta_{d}}$ = 1. The reward $\tilde{r}^{i}_{t}$ of an agent controlling an individual household consists of two components as described in Eq. (4): the negative electricity cost
$R_{t}^{i}=-p_{t}^{i}\,\phi_{t}$
and the safety penalty
$R^{i}_{\textnormal{penalty}}(u^{i}_{t},\tilde{u}_{t}^{i})=-w\|\tilde{u}^{i}_{t}-u^{i}_{t}\|_{2},$
where $w$ is a scaling factor. In the decentralized scenario, each RL controller has a safety shield that guards household-specific constraints.
For the centralized scenario, the agent has access to the observations and actions of all households, however, the electricity prices are only observed once. Since the centralized agent also controls how much electricity has to be imported to balance the system ($p^{M}_{t}$), its action is $u_{t}=\begin{bmatrix}u_{t}^{1}&...&u_{t}^{14}&p^{M}_{t}\end{bmatrix}$. The reward of this agent is the negative cost of the imported electricity
$R_{t}=-\phi_{t}\sum_{i}p_{t}^{i}.$
Since the central agent controls all households within the system, the corresponding safety shield has to ensure the DC power flow constraints in addition to the household constraints. The safety penalty added to the reward is thus
$R_{\textnormal{penalty}}(u_{t},\tilde{u}_{t})=-w\|\tilde{u}_{t}-u_{t}\|_{2}.$
We train the centralized RL agent using the single-agent proximal policy optimization (PPO) [70] implementation by StableBaselines [64].
VI-D Results and Discussion
Because economic dispatch is intrinsically an infinite horizon problem, it is non-trivial to compare finite horizon control performance. Therefore, when evaluating the total household cost, we consider a time period and compare the realized cost from this period plus an approximation of the potential cost of the following day. For this, we compute the system trajectory of the following day, assuming a centralized optimal controller and the available forecasts.
Tab. III summarizes the results for both the centralized and the decentralized control scheme for three different time periods of 10 days. For all time periods and control schemes, the optimal control baseline outperforms the RL controllers for the chosen hyperparameters. Notably, the gap in performance decreases when using forecasts with a uniformly sampled noise. This suggests that the agents trained with the noisy forecasts learn a more robust policy, which generalizes better to unseen data. Furthermore, we can observe that in the decentralized scenario, the RL agents learn to reduce the number of action adjustments. Fig. 5 shows how the number of action adjustments per agent decreases during the course of the training. In the centralized control scheme, the agent learns to reduce the penalty but cannot reduce the number of action adjustments. This is due to the complexity of the constraints: In the centralized scenario, the safety shield has to ensure satisfaction of the DC power flow constraints. This is also observed during deployment: Fig. 6 shows that the actions of the centralized RL agent have to be adjusted in every time step, while the decentralized agent does not require any adjustments for the given day. The results show that RL training is challenging, which is a clear motivation for CommonPower as a platform for developing, studying, and comparing novel strategies and algorithms in a single consistent framework.

### v3

@@ V Control
V-A Unified Control Approach
To enable both the deployment and the training of single-agent and multi-agent systems with heterogeneous controllers, CommonPower offers one unified interface based on the gymnasium API [18]. Developing our own interface was necessary because no standard environment representation has emerged for multi-agent RL thus far. We therefore have an internal environment representation and use wrapper classes to adapt to representations used by the respective RL libraries such as StableBaselines [65] or MAPPO [54].
Our interface realizes the interaction of any control algorithm with the underlying power system. The control flow for one episode with $T$ time intervals is provided by Alg. 1: In each time step $t$, the system loops over all controllers, which compute the control inputs for their respective subsystems. During training, the RL agents will sample these control inputs from their current policies. When training has been completed, each RL agent predicts the control inputs deterministically based on its learnt policy. We refer to this mode as deployment. Subsequently, all control inputs are applied simultaneously to advance the discrete-time system dynamics and transition to the next state.
Algorithm 1 Single run
$t\leftarrow t_{0}$
while $t<t_{0}+T$ do
     $\hat{d}_{[t+1,t+T]}\leftarrow$ forecasts(t)
     $u_{t}\leftarrow$ get_controller_inputs($o_{t}$)
     $c_{t}\leftarrow C(x_{t},d_{t},u_{t})$$\triangleright$ Compute cost
     $x_{t+1}\leftarrow$ additional_updates(t)
     $t\leftarrow t+1$triangleright Step system forward
end while
We want to demonstrate the effectiveness of our unified approach with a small example. The pseudo-code in Fig. 3 shows how to use CommonPower to set up a power system consisting of two buildings. One is controlled by an RL agent, the other by an optimal controller. To the best of our knowledge, no existing tool is able to realize such a setup with heterogeneous controllers. With CommonPower, the user only has to specify the safeguarding mechanism and the RL algorithm. The RL agent can be trained with a single line of code. After training, the system with the two controllers can be simulated using the same interface.
@@ V-A Unified Control Approach
V-B Safeguarding RL-based Controllers
@@
To train and deploy RL-based controllers in safety-critical real-world applications, such as power systems, constraints of the controlled system have to be fulfilled at all times. In the case of decentral control with shared constraints, we adopt a two-stage approach: First, each agent computes its dispatch, and then the grid operator computes the dispatch of balancing assets to fulfill global constraints. This is in line with existing work [38, 39, 32, 42, 41]. Furthermore, it is privacy-preserving since each coalition can compute its own dispatch subject to local constraints based on purely local observations, and the grid operator only needs information about the total energy consumption.
@@
The safeguard of an individual agent solves an optimal control problem that captures the constraints of the subsystem under control over a horizon $T$. The output of the safeguard, the safe action for the current time step $t$, is passed to the environment as shown in Fig. 1. The unique feature of CommonPower is that the constraints of the subsystem are automatically extracted from the symbolic domain of the respective entities as described in Sec. IV-B.
The safeguarding explicitly considers forecast uncertainty, which requires set-based adjustments to the problem formulation from (1). Instead of considering a specific forecast $\hat{d}_{t}^{i}$, we consider the set of forecasted possible realizations $\hat{\mathcal{D}}_{t}^{i}=[\underline{d}_{t}^{i},\overline{d}_{t}^{i}]$. We write an uncertain state as $\mathcal{X}_{t}^{i}$. We can now reformulate the input constraints based on a set of forecasts and an uncertain state:
$u_{t}^{i}=h(x_{t}^{i},\hat{d}_{t}^{i})\in[\underline{u}^{i},\overline{u}^{i}]\Longrightarrow\,u_{t}^{i}\in\mathcal{U}_{t}^{i}(\mathcal{X}_{t}^{i},\hat{\mathcal{D}}_{t}^{i}),$ (2)
with
$\begin{split}\mathcal{U}_{t}^{i}(\mathcal{X}_{t}^{i},\hat{\mathcal{D}}_{t}^{i})\subseteq\bigg{\{}&u\in[\underline{u}^{i},\overline{u}^{i}]\,\Big{|}\,\\ &\exists x\in\mathcal{X}_{t}^{i},\exists d\in\hat{\mathcal{D}}_{t}^{i}\colon u=h(x,d)\bigg{\}}.\end{split}$ (3)
For convenience, we will denote the feasible input set as $\mathcal{U}_{t}^{i}$ in the remainder of this section. Note that the specific computation of the feasible input set depends on the input coupling $h(\cdot)$. If $h(\cdot)$ is linear, we can directly compute $\mathcal{U}_{t}^{i}$ with any set representation that is closed under the operations of linear map and Minkowski sum222 Given two sets $\mathcal{A},\mathcal{B}\in\mathbb{R}^{n}$ and a matrix $M\in\mathbb{R}^{m\times n}$, we define the linear map as $M\mathcal{A}=\{Ma\mid a\in\mathcal{A}\}$, and the Minkowski sum as ${\mathcal{A}\oplus\mathcal{B}=\{a+b\mid a\in\mathcal{A},b\in\mathcal{B}\}}$. . If $h(\cdot)$ is nonlinear, we can find convex overapproximations, as demonstrated, for example, in [66].
Depending on the realization of forecasted values, we might have to apply any input from $\mathcal{U}_{t}^{i}$. This means we must consider the entire feasible input set in the system dynamics $f(\cdot)$. To this end, we reformulate the difference equation from (LABEL:oc_high_level_coalition_constraints) into a difference inclusion:
$x_{t+1}^{i}=f^{i}(x_{t}^{i},\hat{d}_{t}^{i},u_{t}^{i})\Longrightarrow\,\mathcal{X}_{t+1}^{i}\subseteq f^{i}(\mathcal{X}_{t}^{i},\hat{\mathcal{D}}_{t}^{i},\mathcal{U}_{t}^{i}).$ (4)
Again, the specific set-based computation depends on $f(\cdot)$.
Our implementation in CommonPower exclusively uses intervals as set representations to maximize tractability and requires $f(\cdot)$ to be continuous, injective, and monotonous. Under these conditions, all constraints can be efficiently posed as constraints on the intervals’ upper and lower limits. Note that these conditions are usually not restrictive in practice. CommonPower expects instances of Power Flow entities to return $\hat{\mathcal{D}}_{t}^{i}$ as an interval set. All other functionals are given by the user in the non-set-based form of (1) and automatically converted to their set-based equivalent.
Our safeguarding framework is based on the predictive safety filter from [57] and abstracts all post-posed mechanisms. Specifically, the safeguard finds a safe action $\tilde{u}^{i}_{t}$ based on the proposed action $u^{i}_{t}$ of the RL agent by solving
$\begin{split}&\tilde{u}^{i}_{t}=\operatorname*{argmin}_{u^{\prime}_{t}}\phi(u^{\prime}_{t},u^{i}_{t})\\ &s.t.\\ &u_{t}^{i}\in\mathcal{U}_{t}^{i}\\ &\mathcal{X}_{t}^{i}\subseteq[\underline{x}^{i},\bar{x}^{i}]\\ &\mathcal{X}_{t+1}^{i}\subseteq f^{i}(\mathcal{X}_{t}^{i},\hat{\mathcal{D}}_{t}^{i},\mathcal{U}_{t}^{i}).\end{split}$ (5)
where $\phi(u^{\prime}_{t},u^{i}_{t})$ is a cost function defining the type of mechanism. With $\phi(u^{\prime}_{t},u^{i}_{t})=\|u^{\prime}_{t}-u^{i}_{t}\|_{2}$, the approach is commonly called action projection. In contrast, with $\phi(u^{\prime}_{t},u^{i}_{t})=0$, we obtain random actions with the solver initialization as the source of randomness. This is an example of so-called action replacement [14].
@@
such that the tuple used for learning is $(o^{i}_{t},u^{i}_{t},o^{i}_{t+1},\tilde{r}^{i}_{t})$. This penalty informs the agent about the intervention of the safeguard. How to compute the penalty is a design choice and can require careful tuning.
VI Experiments
Figure 4: Mean rollout length during training for four selected combinations of safeguarding approach and penalty. We provide the mean and standard deviation over five random seeds for each configuration. One rollout comprises three consecutive days, meaning that the maximum rollout length is $72$.
To demonstrate the capabilities of CommonPower, we present a benchmark that compares RL agents trained without safeguarding and with different types of safeguards against an optimal control baseline. The scenario is an energy management system for a battery that optimizes the self-consumption of electricity generated by a photovoltaic module. We employ perfect forecasts, which avoids the influence of forecast uncertainty and enables us to highlight differences in performance between the safeguarding approaches. PPO is used as an RL algorithm across all experiments.
@@ VI-A System Model
Figure 5: State of charge of the battery over one week, starting 30 July 2016. We compare three configurations: The optimal controller, an agent trained without safeguard (but deployed with the projection safeguard), and an agent trained with the projection safeguard.
We model a household with an inflexible load, a battery storage system, and a photovoltaic generator. The task of the controller is to optimize the battery usage in order to minimize electricity costs for the household. The buying price for one kWh is set to $0.37$€ and the selling price to $0.08$€. For the household load data as well as the renewable generation, we use time series provided by the SimBench [67] benchmark data set.
VI-B RL Training
In each time step $t$, the RL agent gets an observation
$o_{t}~{}=~{}\begin{bmatrix}soc_{t}&\hat{\ell}_{[t,T]}^{i}&\hat{g}_{[t,T]}&\hat{\varphi}_{[t,T]}\end{bmatrix}^{T}$
@@ VI-B RL Training
where $soc_{t}$ is the state of charge of the battery and ^$\hat{\ell}_{[t,T]}$, $\hat{g}_{[t,T]}$, and $\hat{\varphi}_{[t,T]}$ are the forecasted load, generation, and electricity prices. We choose a forecast horizon of 12 time steps and a time interval of one hour. The action of the agent is the charging/discharging power of the battery, $u_{t}=\begin{bmatrix}p_{t}^{B}\end{bmatrix}.$ The reward $\tilde{r}_{t}$ consists of two components as described in Eq. (6): the negative electricity cost $R_{t}=-p_{t}\,\varphi_{t}$ and a penalty $R_{\textnormal{penalty}}$.
When training without safeguarding, we terminate the episode and reset the system as soon as a constraint – here the maximum or minimum battery capacity – is violated. We evaluate two different types of safeguards:
  • action projection, where the safeguard finds the safe action that is closest to the one proposed by the RL agent, and
  • action replacement, where a random safe action is chosen.
Each safeguarding option is combined with different penalty options. We either add
  • no penalty or
  • a constant penalty of magnitude $R_{\textnormal{penalty}}=-1.0$
to the reward. In addition, we experiment with a distance depending penalty. When using action projection safeguarding, this penalty is proportional to the distance between the safe and the unsafe action, such that $R_{\textnormal{penalty}}=-w\|\tilde{u}_{t}-u_{t}\|_{2},$ where $w$ is a scaling factor. We did not investigate this option for the replacement safeguard as it does not provide any useful information to the agent. For training without safeguarding, the penalty is proportional to the percentage of completed steps before a constraint violation occurs.
TABLE III: Deployment results over 5 training seeds and 5 evaluation periods, each one week long. We report the cumulative cost over the five weeks as a mean over the training seeds for the RL controllers. The mean number of interventions refers to how often the safeguard had to intervene during one evaluation period to prevent constraint violations.
Costs Interventions
Optimal controller 18.55 0.0
RL controller trained without safeguard no penalty 24.58 2.44
constant penalty 22.78 2.96
dd penalty 26.08 2.2
RL controller trained with projection safeguard no penalty 24.50 2.32
constant penalty 26.96 0.08
dd penalty 25.47 3.2
RL controller trained with replacement safeguard no penalty 27.05 0.4
constant penalty 26.57 0.12
dd penalty n/a n/a
VI-C Results and Discussion
Each configuration of RL agents is trained over five random seeds. Fig. 4 visualizes the average rollout length during training for some selected combinations of safeguard and penalty. When training without a safeguard, constraint violations occur frequently and induce a reset of the environment. Therefore, the mean rollout length is shorter. Fig. 4 illustrates that the rollout length can be increased by using a penalty whenever we have to reset.
During deployment, we no longer allow constraint violations and add an action projection safeguard to all RL agents trained without safeguarding. While the state of charge of the battery was randomly initialized during training, it is fixed to half of the capacity during deployment. Tab. III reports the sum of the electricity costs over five evaluation periods, starting February 1, April 15, July 30, September 23, and November 10, and lasting one week. Furthermore, Tab. III displays the mean number of interventions during one evaluation period, with the theoretical maximum being $7*24=168$ interventions. The results are averaged over the five training seeds.
The best performance can be achieved with the RL agents trained without a safeguard and with a constant penalty. Fig. 5 shows that an agent with this configuration has learned a slightly more varying behavior compared to an agent trained with a projection safeguard and a constant penalty. The latter exhibits a very similar charging and discharging behavior for every day, regardless of the actual amount of surplus energy. We hypothesize that the agent trained without a safeguard achieved better exploration because its behavior policy was the same as the policy being optimized during training. This is a central paradigm of on-policy algorithms such as PPO, and adding a safeguard during training changes it. Further experiments are required to establish whether this phenomenon prevails. In any case, a safeguard is necessary whenever interaction with a real system is required during training.
Furthermore, Tab. III shows that using a constant penalty when training with a safeguard can decrease the number of safeguard interventions during deployment. This can induce a rise in costs, possibly because the maximum and minimum capacity of the battery are exploited less. However, obtaining an agent that relies as little as possible on the safeguard can be desirable, for example, when real-time requirements exist and safeguarding involves time-consuming computations.
The experiments demonstrate a complex interaction between the safeguards, the safety penalty, and the ability of the agent to learn cost optimization versus minimization of safety interventions. Investigations along more dimensions, such as further single-agent RL algorithms, multi-agent RL, or imperfect forecasts, are required to enable robust statements on the performance of RL under different safeguarding approaches. With CommonPower, such studies can readily and consistently be conducted across a variety of scenarios.